Let D(p) = 4-p and S(p) = 1 + p. Using the method of linear first-order differential equations, find a general solution to p(t) (it will involve k). What is the long term behaviour of the price? Does it tend to a specific value regardless of the initial price?
Answer #1

Q). Ot = 4-Pt St= It Pet, at Eqm Dz= St 4-Pt = 1HPt- 3-Pt a = Pt. 80 fiast order difference En. I PE = 3-Pt-1 Comparing to st

thus Price doesn't tend to specific value , regardless of intial value, rather it oscillates